OpenStudy (annipuppi):

One quick geometry question. I have the answer, but I need the work. It's a test correction and I don't understand how to solve it.

OpenStudy (annipuppi):

The answer is 27

OpenStudy (annipuppi):

@surjithayer @ganeshie8 please, I really need help on this. I've tried so many ways of solving it, but I can't come up with the right answer... how do I do it?

ganeshie8 (ganeshie8):

http://www.regentsprep.org/Regents/math/geometry/GP15/CircleAngles.htm scroll down to the last

OpenStudy (annipuppi):

THANK YOU! so u=1/2(120-66) u=1/2(54) u=27?

ganeshie8 (ganeshie8):

since the inscribed angle \(\angle S = 33\), near arc = \(33\times 2 = 66\) apply the formula : \(\angle U = \dfrac{\text{(far arc) - (near arc)}}{2} = \dfrac{120-66}{2} = 27\)

OpenStudy (annipuppi):

Okay, thank you so much!

ganeshie8 (ganeshie8):

yw